decompose "g(x)=5\/(x-1)^3"


1
Expert's answer
2022-02-07T15:53:07-0500

We are looking for two functions, f and h such that

"g(x)=f(h(x))"

Here, "g(x)=\frac{5}{(x-1)^3}"

We can take "f(x)=\frac{5}{x^3}\ \text{and}\ h(x)=(x-1)" such that "g(x)=f(h(x))""=\frac{5}{(x-1)^3}"

We could then decompose the function as

"f(x)=\frac{5}{x^3}\ \text{and}\ h(x)=(x-1)"